OpenStudy (joshawad):

anybody know anything about Solving Simultaneous equations By Substitution I know I'm supposed to pick numbers an d put them into my equation for a graph (in my case -6 to +6 but I still have 1 question how do I know when I have the right number

OpenStudy (anonymous):

no sorry im to blonde :( good luck thoigh

OpenStudy (joshawad):

then don't respond...

OpenStudy (mathmale):

Hey, Josh, how about posting those simult. eqns.?

OpenStudy (joshawad):

Line 1: Y - X = 0 Line 2: Y - 3X - 4

OpenStudy (mathmale):

Josh, if you're to solve this system using the subst. method, then please just solve one of the equations for either x or y. I'd do that to the first equation, because it's simpler. Show me your stuff.

OpenStudy (joshawad):

well what ever X is in line 1 Y is the opposite

OpenStudy (joshawad):

so prob start on line 2:

OpenStudy (mathmale):

You could start with either one! I've suggested starting with the first equation because there's less work involved there. Your task is to solve the first equation for y. Here's what I'd do: Write the equation as it is now: y-x=0 Add x to both sides: y - x = 0 +x = +x ------------- Could you do that?

OpenStudy (joshawad):

y=0 but that's not possible

OpenStudy (mathmale):

Josh: Here's what I'd get: y-x=0 x x -------- y = x Done. Be sure you're comfortable with this before we move on.

OpenStudy (joshawad):

but if u just got rid of x....

OpenStudy (mathmale):

What I did was to move that x from the left side to the right side of the equation. The new version of this equation tells us exactly the same information as does the original one. "If equals are subtracted from equals, the resulting terms are equal." For now, please accept the fact that y=x.

OpenStudy (joshawad):

OH I see it

OpenStudy (mathmale):

Now we act on the fact that y=x. In the second equation, y-3x=4 (or at least, I'm assuming that that's what you meant). We substitute y=x. Then x-3x=4. Can y ou solve that for x? Hint: combine x and =3x first.

OpenStudy (mathmale):

How are you doing?

OpenStudy (joshawad):

ao x is 2

OpenStudy (mathmale):

Josh, we have this equation: x-3x=4 and we want to solve this equation for x. Combining x and -3x results in -2x, right? then -2x=4 Divide both sides of this equation by -2 to obtain a value for x. (It's not 2).

OpenStudy (joshawad):

X=2 Y=2

OpenStudy (mathmale):

If you and I get different results, it's probably because your second equation was typed as y-3x-4. Actually, Josh, that is NOT an equation. Please go back to the orginal problem and type what you meant in the 2nd line.

OpenStudy (mathmale):

Happy to "meet" and work with you, Josh. Unfortunately, I need to get off my computer. Try posting a similar problem; perhaps someone else can help you apply the same principles to solving this new system of equations. Best of luck to you! MM

OpenStudy (anonymous):

y - 3x = 4 y = 3x + 4 now sub 3x + 4 in for y in the other equation y - x = 0 3x + 4 - x = 0 2x + 4 = 0 2x = -4 x = -2 now sub -2 in for x in either of the equations to find y y - 3x = 4 y - 3(-2) = 4 y + 6 = 4 y = 4 - 6 y = -2 check... y - x = 0 y - 3x = 4 -2 - (-2) = 0 -2 - 3(-2) = 4 -2 + 2 = 0 (correct) -2 + 6 = 4 4 = 4 (correct) your solution is (-2,-2)

OpenStudy (joshawad):

ok what about 2x-y=4 y=-x+11

OpenStudy (joshawad):

main ly 2x-y=4 and lol

OpenStudy (joshawad):

@kelliegirl33

OpenStudy (joshawad):

@mjbmanny

OpenStudy (anonymous):

y = -x + 11 now sub -x + 11 in for y in the other equation 2x - y = 4 2x - (-x + 11) = 4 2x + x - 11 = 4 --- add 11 to both sides 2x + x = 4 + 11 3x = 15 x = 5 now sub 5 in for x into either of your equations to find y y = -x + 11 y = -5 + 11 y = 6 check.. 2x - y = 4 2(5) - 6 = 4 10 - 6 = 4 (correct) solution is (5,6)

OpenStudy (joshawad):

how did u get x=-11 in line 5 @kelliegirl33

OpenStudy (anonymous):

2x - (-x + 11) = 4 --- distribute the - through the parenthesis. If it makes it easier, if you just have a sign, with no number, you can put a 1. Like this... 2x - 1(-x + 11) = 4 -- now distribute through the parenthesis 2x + x - 11 = 4 Understand ?

OpenStudy (joshawad):

no

OpenStudy (anonymous):

do you know how to distribute through the parenthesis ?

OpenStudy (joshawad):

the distributive property yes

OpenStudy (anonymous):

thats all I did. 2x - (-x + 11) = 4 If you distributed the negative sign through the parenthesis, what would you get ?

OpenStudy (anonymous):

or write it like this... 2x - 1(-x + 11) = 4

OpenStudy (joshawad):

2x + -x -11 =4

OpenStudy (anonymous):

negative times a negative is a positive - * -x = x....not -x

OpenStudy (joshawad):

oh right

OpenStudy (joshawad):

y= -5+11 y=6?

OpenStudy (anonymous):

so that leaves : 2x + x - 11 = 4 any more questions ?

OpenStudy (joshawad):

correct?

OpenStudy (anonymous):

y = -5 + 11 when you are adding/subtracting different signs, you subtract and take the sign of the bigger number. y = 6

OpenStudy (anonymous):

correct..y = 6

OpenStudy (joshawad):

(5,6)

OpenStudy (anonymous):

correct

OpenStudy (joshawad):

now my probem is i have to draw the line remember how i said earlier that my grah\ph only goes to -6 to +6

OpenStudy (anonymous):

you can always check your answers by subbing them into either/both of you equations. If they come out equal, they are correct

OpenStudy (joshawad):

never mind lol it said it was bigger

OpenStudy (joshawad):

but what ot should be at it still confusing

OpenStudy (anonymous):

what ?

OpenStudy (joshawad):

i havr to draw the line still

OpenStudy (anonymous):

what equation are you trying to graph ?

OpenStudy (joshawad):

both

OpenStudy (joshawad):

that's what the point is for

OpenStudy (anonymous):

you are trying to graph y - x = 0 and y - 3x = 4 ??

OpenStudy (joshawad):

never mind that 2x +Y=-1 ho do i change that without - -1 or should it be - -1

OpenStudy (anonymous):

I do not understand what you are trying to do. Are you trying to put it in y = mx + b form ?

OpenStudy (joshawad):

equation of the line like y=-3x+8

OpenStudy (anonymous):

you want me to show you how to graph y = -3x + 8 ??

OpenStudy (joshawad):

NO! omg.... 2x+y=-1 i made that into 2x=y- -1 is tha6 correct?

OpenStudy (anonymous):

2x + y = -1....why are you changing it to 2x = -y - 1 ? It would be easier to solve for y, rather then x.

OpenStudy (joshawad):

...

OpenStudy (joshawad):

ur missing the point

OpenStudy (anonymous):

2x + y = -1 2x = -y - 1 x = -y/2 - 1/2 or x = -1/2y - 1/2

OpenStudy (joshawad):

2x + y = -1 is easier to graph? no i don't think you can graph that

OpenStudy (anonymous):

to graph an equation, you need to put it in y = mx + b form

OpenStudy (joshawad):

i know that!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

2x + y = -1 y = -2x - 1 y intercept is (0,-1) x intercept can be found by subbing in 0 for y : 2x + 0 = 1 2x = 1 x = 1/2....x intercept is (1/2,0) slope is -2

OpenStudy (joshawad):

what

OpenStudy (joshawad):

2x + y = -1to 2x = y - -1 is that right yes or no

OpenStudy (anonymous):

no 2x + y = -1 ----- subtract y from both sides 2x = -y - 1

OpenStudy (joshawad):

what happened to the negative on the 1

OpenStudy (anonymous):

it stayed the same...we did not even mess with the -1

OpenStudy (joshawad):

yes u did

OpenStudy (joshawad):

1 had a negative on it!!!!!!!

OpenStudy (anonymous):

2x + y = -1 --- subtract y from both sides 2x + y - y = -y - 1 -- combine like terms 2x = -y - 1

OpenStudy (joshawad):

THE NEGATIVE!

OpenStudy (joshawad):

on the 1

OpenStudy (anonymous):

what about the negative on the 1 ? It is still there.

OpenStudy (joshawad):

then how about the sign sitting therei wont nibpick at the space u put in there but there is no sign between the -y and the -1

OpenStudy (anonymous):

there does not have to be a sign.....2x = -y - 1 is correct. But if you want to isolate x, you need to divide everything by 2

OpenStudy (anonymous):

it is not supposed to be 2x - y - (-1) because 2x+ y = -1 ---- subtract y from both sides 2x = -y - 1

OpenStudy (anonymous):

do you understand ?

OpenStudy (joshawad):

i think so working on a problem

OpenStudy (anonymous):

ok....tag me if you need help....that is, if you want to :)

OpenStudy (joshawad):

ok i will

OpenStudy (joshawad):

@kelliegirl33 2X+ Y = -1 and Y=-3X take -3X and sub for Y 2X + (-3X) = -1 add -3X 2X+3X= -1 5x =-1 what did i do wrong?

OpenStudy (joshawad):

@kelliegirl33

OpenStudy (anonymous):

2x + y = -1 y = -3x sub 2x +(-3x) = -1 2x - 3x = -1 -x = -1 x = 1

OpenStudy (anonymous):

you did not have to add -3x. All you had to do was distribute the + through the parenthesis. This gave us 2x - 3x = -1. We then combined like terms. -x = -1. I then multiplied by -1 to make x positive.

OpenStudy (joshawad):

oh

OpenStudy (anonymous):

understand ?

OpenStudy (joshawad):

:o

OpenStudy (anonymous):

believe me, when you do enough of these, you will get it

OpenStudy (joshawad):

this geometry/ algebra III gene doesn't run in my family

OpenStudy (anonymous):

geometry...yuck...I am so not good at geometry. Algebra is ok. It just takes a lot of practice to learn all this. I will probably not even need this in my future. Just learn it to get out of high school.....lol

OpenStudy (joshawad):

hate to have to break it to you but this is a crossover of geometery

OpenStudy (joshawad):

:)(weak smile)

OpenStudy (joshawad):

YES OMG YES!!!!!!!